OpenStudy (jenniferjuice):

WHAZZZZZAAAA HALP Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below birth month: 8 birthday: 16

OpenStudy (anonymous):

I don't think you need help with this.

OpenStudy (anonymous):

What's the f(x) equation in vertex form?

OpenStudy (jenniferjuice):

I DONT KNOW :(

OpenStudy (anonymous):

Lol ok, y = a(x-h)² + k a= 4 - (-4) h=8 k=16 so, y=4(x-8)²+16

OpenStudy (anonymous):

that's all you had to do ;)

OpenStudy (jenniferjuice):

well thanks you :D medal for you! B)

OpenStudy (anonymous):

aww... thank you. good luck on the rest